# Taxonomizing Representational Harms using Speech Act Theory

### v1

@@ Abstract
Representational harms are widely recognized among fairness-related harms caused by generative language systems. However, their definitions are commonly under-specified. We present a framework, grounded in speech act theory (Austin, 1962), that conceptualizes representational harms caused by generative language systems as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). Building on this argument and drawing on relevant literature from linguistic anthropology and sociolinguistics, we provide new definitions stereotyping, demeaning, and erasure. We then use our framework to develop a granular taxonomy of illocutionary acts that cause representational harms, going beyond the high-level taxonomies presented in previous work. We also discuss the ways that our framework and taxonomy can support the development of valid measurement instruments. Finally, we demonstrate the utility of our framework and taxonomy via a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured.
@@ 1 Introduction
Representational harms (Barocas et al., 2017; Crawford, 2017)—i.e., “[harms that] arise when a system represents some social groups in a less favorable light than others, demeans them, or fails to recognize their existence altogether” (Blodgett, 2021)—are widely recognized among fairness-related harms caused by generative language systems, including LLM-based systems, even though definitions of these harms are commonly under-specified (Blodgett et al., 2020; Rauh et al., 2022). This under-specification is further compounded by the way “harms” are typically discussed in the AI literature, referring sometimes to types of system behaviors, sometimes to the impacts of those system behaviors, and sometimes to broader societal impacts arising from the development and deployment of AI systems (e.g., Banko et al., 2020; Weidinger et al., 2022; Shelby et al., 2023; Solaiman et al., 2023; Abercrombie et al., 2024; Hutiri et al., 2024; Slattery et al., 2024; Zeng et al., 2024).
This lack of conceptual clarity makes developing valid measurement instruments and effective mitigations fraught. For example, it makes it difficult to understand the precise concept(s) that a given measurement instrument or mitigation is targeting. And, in the case of measurement instruments, it makes it difficult to understand what the resulting measurements do and do not mean. To address this challenge, we present a framework grounded in speech act theory (Austin, 1962), shown at a high level in Figure 1 and described in Section 2, that conceptualizes representational harms as the perlocutionary effects, (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). This framework enables us to draw clearer distinctions between system behaviors and their impacts, and to provide new definitions of stereotyping, demeaning, and erasure. In Section 3, we then use our framework to develop a granular taxonomy that highlights distinguishing features of the types of illocutionary acts—i.e., system behaviors—that cause representational harms.
@@
Figure 1: A visualization of our framework of representational harms, described in detail in Section 2.
@@ 2 A Speech Act Theoretical Framework of Representational Harms
Representational harms are widely recognized among fairness-related harms caused generative language systems (e.g., Barocas et al., 2017; Crawford, 2017; Blodgett, 2021; Katzman et al., 2023). However, they are often presented as high-level concepts—specifically, stereotyping, demeaning, and erasure—with corresponding definitions that lack internal theoretical coherence and do not provide the granularity needed to develop valid measurement instruments and effective mitigations. In this section, we explain how speech act theory (Austin, 1962) provides a coherent conceptual framework for representational harms.
@@ 2.1 Representational Harms as Perlocutionary Effects
Building on the argument above, we conceptualize all harms caused by generative language systems—including representational harms—as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). Specifically, representational harms occur when the perlocutionary effects of a system output include the entrenchment—i.e., the further cementing in the world—of one or more harmful [^1] One could argue that all social hierarchies are inherently harmful because they differentially confer power, status, privileges, resources and opportunities. That said, we intentionally include the word “harmful” to emphasize our focus on harms. For a broad range of alternative perspectives on social hierarchies and power, see Bourdieu (1984), Connell and Messerschmidt (2005), Gramsci (1971), Sidanius and Pratto (1999), Diberardino et al. (2024), and Anderson and Brown (2010). social hierarchies, where a social hierarchy is a systematic organizations of individuals or social groups that differentially confers power, status, privileges, resources, and opportunities. We provide more details about individuals and social groups within harmful social hierarchies and how they relate to questions about identity formation and maintenance in Appendix B.
Harmful social hierarchies are entrenched when representations of the world that involve those hierarchies are (re-)produced—e.g., in the outputs of generative language systems. As a result, the entrenchment of one or more harmful social hierarchies constitutes a first-order perlocutionary effect because it occurs at the time of output generation. However, representations of the world that involve harmful social hierarchies can also influence individuals’ beliefs and psychological states—e.g., by causing someone to feel harmed. These real-world impacts can be either first-order or second-order perlocutionary effects because they can occur at the time of output generation or subsequent to it. We provide additional theoretical considerations relating to the entrenchment of harmful social hierarchies in Appendix B. We also explain how other fairness-related harms, namely allocation harms and quality-of-service-harms, can be similarly conceptualized as the perlocutionary effects of particular types of illocutionary acts in Appendix D.
@@ 2.2 Stereotyping, Demeaning, and Erasure as Types of Illocutionary Acts
We conceptualize stereotyping, demeaning, and erasure as particular types of illocutionary acts, or system behaviors, whose perlocutionary effects include the entrenchment one or more harmful social hierarchies. Collectively, these types of illocutionary acts span the five basic classes of illocutionary acts proposed by Searle (1976).
Building on existing high-level definitions of stereotyping, demeaning, and erasure that relate these concepts to normative views about identity (Blodgett, 2021), we draw on speech act theory, linguistic anthropology, and sociolinguistics to provide new definitions of these concepts. Specifically, our definitions rely on the linguistic anthropological notion of evaluative lenses that can be be used to either empower or disempower individuals and social groups when using language to 1) characterize social groups, 2) position individuals within social groups, and 3) position social groups within harmful social hierarchies. In other words, these evaluative lenses can be used to either disrupt or entrench harmful social hierarchies. We consider two such evaluative lenses: similarity/difference and authenticity/inauthenticity. [^4] We adapted these evaluative lenses from a set of semiotic and social processes—called the tactics of intersubjectivity—that are identified by linguistic anthropologists as being critical to identity formation. For a review of these processes, see Bucholtz and Hall (2004) and Bucholtz and Hall (2005)’s discussions of adequation and distinction, authentication and denaturalization, and authorization and illegitimation. We provide a detailed discussion of these lenses in Appendix E.
We provide our new definitions of stereotying, demeaning, and erasure below. We illustrate these definitions using example utterances from ToxiGen Hartvigsen et al. (2022), a widely used, large-scale, machine-generated dataset for training classifiers to measure implicit toxicity. We focus specifically on three utterances that target LGBTQ+ people and provide a table summarizing their types of illocutionary acts, illocutionary act patterns, evaluative lenses, and illocutionary effects in Appendix H.
@@ 2.2.1 Stereotyping
As a type of behavior, stereotyping has previously been characterized as involving stereotyping beliefs—i.e., “fixed, over generalized belief[s] about […] particular group[s] of people” (Cardwell, 1996). [^5] Some definitions of stereotyping involve positive or negative sentiments. In contrast, we take the view that stereotyping spans positive, negative, and neutral sentiments. Building on this definition, we define a stereotyping illocutionary act to be an illocutionary act that 1) invokes both within-group similarity and between-group difference and 2) implicitly positions the referenced characteristic(s) as indicator(s) of authenticity, reinforcing the criteria by which individuals are positioned within or excluded from the targeted social group(s). These evaluative lenses are used to empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)). Specifically, a stereotyping illocutionary act essentializes the members of each targeted social group by characterizing them as being similar to one another and different from members of one or more other social groups on the basis of one or more characteristics. For example, the utterance “Gay men are not very masculine” essentializes gay men by characterizing them as being as similar to one another and, implicitly, different from straight men—on the basis of being not very masculine. The characteristic of being not very masculine is also positioned as an indicator of authenticity for gay men. We note that stereotyping illocutionary acts are typically representatives Searle (1976).
@@ 2.2.2 Demeaning
Demeaning has previously been characterized as involving “dehumanizing or offensive associations, or [behaviors that] otherwise threaten people’s sense of security or dignity” (Blodgett, 2021). We build on this definition and define a demeaning illocutionary act to be an illocutionary act that disempowers one or more social groups (or one or more individuals based on their membership in those social group(s)) by invoking between-group difference and, often implicitly, within-group similarity. In contrast to a stereotyping illocutionary act, a demeaning illocutionary act does not invoke the evaluative lens of authenticity/inauthenticity. In other words, while a stereotyping illocutionary act positions the referenced characteristic(s) as indicator(s) of authenticity, a demeaning illocutionary act does not. For example, the utterance “Obviously bisexuals are messed up” disempowers bisexual people by stigmatizing them on the basis of deviance. This utterance characterizes bisexual people as being different from non-bisexual people without positioning any characteristic(s) as indicator(s) of authenticity. Unlike stereotyping illocutionary acts, which are typically representatives, demeaning illocutionary acts can be representatives, expressives, directives, or commissives Searle (1976).
@@ 2.2.3 Erasure
Finally, erasure has previously been characterized as a failure to recognize the existence of social groups, often by “foregrounding dominant understandings and […] ideologies” (Blodgett, 2021). [^6] We note that this is a contextually specific definition that aligns with established theoretical work on erasure in linguistic anthropology (see, e.g., the work of Irvine and Gal, 2000). We define an erasing illocutionary act to be an illocutionary act that invokes within-hierarchy similarity to disempower one or more social groups (or one or more individuals based on their membership those social group(s)). An erasing illocutionary act erases differences within a harmful social hierarchy by characterizing that hierarchy as being simpler or more internally similar than it actually is, or by characterizing the members of multiple social groups as being similar to one another on the basis of one or more characteristics. In other words, an erasing illocutionary act fails to recognize socially meaningful differences within a harmful social hierarchy. For example, the utterance “There is no way that bisexuality is a real thing” presents the hierarchy of possible sexualities as simpler—e.g., comprised of a smaller set of social groups, or involving fewer characteristics—than it actually is. By invoking within-hierarchy similarity instead of between-group similarly, this definition captures an expansive understanding of erasure that highlights the role of harmful social hierarchies themselves. Erasing illocutionary acts can be representatives, directives, commissives, or declarations (Searle, 1976).
@@ 3 A Granular Taxonomy of Stereotyping, Demeaning, and Erasure
Stereotyping Describes stereotypically "All [social group] are [characteristic]" Group/individual is stereotyped
@@
Caricatures individual as a member of a social group "Those Russian trolls are always at it"
Prescribes stereotypically "[Social group] should be [characteristic]"
Proscribes stereotypically "[Social group] shouldn’t be [characteristic]."
@@
Demeaning Stigmatizes using slurs "R*tard!" Group/individual is stigmatized
@@
Stigmatizes using general insult "Those [social group] are such jerks"
@@
Stigmatizes on basis of disorder, disease, uncleanliness, or untidiness "[Social group] are riddled with disease"
@@
Stigmatizes on basis of immorality, danger, or deviance "Being [social group] is a sin"
@@
Stigmatizes on basis of lack of social capital (popularity, likability, interest, success, or other positive quality) "[Social group] are not that successful, are they?"
@@
Advocates for stigmatizing treatment/avoidance "Avoid [social group] at all costs!"
@@
Threatens or promises stigmatizing treatment/avoidance "Better keep some distance from [social group] or you’ll end up like them"
@@
Others individuals from group as poor exemplar "They’re not a real [social group member]" Group/individual is othered
@@
Others group from (dominant) group "[Social group] are just really different from the rest of us"
@@
Others subgroup from superset group "[Social subgroup] aren’t really even [social superset group]"
@@
Advocates for treatment as other "Let’s make sure [social group] can’t get too comfortable here"
@@
Threatens or promises othering treatment "If they’re not careful, [social group] will end up being treated differently from everyone else"
@@
Objectifies (reduces to physical attribute or behavior) "[Social group] are all legs!" Group/individual is simplified
@@
Glorifies (reduces to prominent and often flattering character or personality traits) "[Social group] are so brave in the face of so much adversity"
@@
Advocates for treatment along the lines of simplified (objectifying or glorifying) features "Let’s go get us some [social group] ass"
@@
Threatens or promises treatment along the lines of simplified (objectifying or glorifying) features "We’re going to treat you [social group] like your [objectified body parts] deserve"
@@
Equates to baby/child, animal, disease or inanimate object "[Social group] are such babies Group/individual is dehumanized
@@
Notes characteristics that match baby/child, animal, disease or inanimate object "Why are [social group] always whining?"
@@
Advocates for treatment like baby/child, animal, disease or inanimate object "[Social group] need to be disciplined"
@@
Threatens or promises treatment like baby/child, animal, disease or inanimate object "If those [social group] aren’t careful, I’m going to skewer them"
@@
Erasing Deprioritizes or questions the needs that diverge from the norm or majority "There aren’t that many [social group] here anyway, so consideration for their needs isn’t important" Group’s/individual’s needs, experiences, contributions and accountability are presented as equal
@@
Delegitimizes or denies a group’s suffering or minimizes the harm of oppression, including via denial of historical events "[Social group persecution event] wasn’t that bad for [social group] anyway"
@@
Diffuses accountability for a harm across the hierarchy, including blaming victims for their circumstances "[Social group] shouldn’t have been [characteristic] if they didn’t want that to happen to them"
@@
Advocates for non-differentiated treatment "[Social group] shouldn’t get special treatment just because they’re [characteristic]"
@@
Threatens or promises non-differentiated treatment "I guarantee we’re not going to specially accommodate [social group]"
@@
Denies justice and fair access to accommodations "No exceptions for [social group]!"
@@
Conflates individuals or social groups "Aren’t [social group 1] and [social group 2] basically the same?" Groups/individuals are homogenized and presented as indistinguishable
@@
Denies existence or fails to recognize individuals or groups "There’s no such thing as [social group]"
@@
Denies existence of individual social group members with certain characteristics "I’ve never met any [characteristic] [social group]"
Advocates for exclusion "Don’t let any [social group] in!"

### v2

@@ Abstract
Representational harms are widely recognized among fairness-related harms caused by generative language systems. However, their definitions are commonly under-specified. We make a theoretical contribution to the specification of representational harms by introducing a framework, grounded in speech act theory (Austin, 1962), that conceptualizes representational harms caused by generative language systems as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). Building on this argument and drawing on relevant literature from linguistic anthropology and sociolinguistics, we provide new definitions of stereotyping, demeaning, and erasure. We then use our framework to develop a granular taxonomy of illocutionary acts that cause representational harms, going beyond the high-level taxonomies presented in previous work. We also discuss the ways that our framework and taxonomy can support the development of valid measurement instruments. Finally, we demonstrate the utility of our framework and taxonomy via a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured.
@@ 1 Introduction
Representational harms (Barocas et al., 2017; Crawford, 2017)—i.e., “[harms that] arise when a system represents some social groups in a less favorable light than others, demeans them, or fails to recognize their existence altogether” (Blodgett, 2021)—are widely recognized among fairness-related harms caused by generative language systems, including LLM-based systems, even though definitions of these harms are commonly under-specified, leading to conceptual confusion, invalid measurement instruments, and ineffective mitigations (Blodgett et al., 2021; Blodgett, 2021; Wallach et al., 2025). This under-specification is further compounded by the way “harms” are typically discussed in the AI literature, referring sometimes to types of system behaviors, sometimes to the impacts of those system behaviors, and sometimes to other broader societal impacts arising from either the development or the deployment of these generative language systems (e.g., Banko et al., 2020; Weidinger et al., 2022; Shelby et al., 2023; Abercrombie et al., 2024; Hutiri et al., 2024; Slattery et al., 2024; Zeng et al., 2024).
This lack of conceptual clarity makes developing valid measurement instruments and effective mitigations fraught. For example, it makes it difficult to understand the precise concept(s) that a given measurement instrument or mitigation is targeting (Blodgett et al., 2020, 2021; Dev et al., 2022). And, in the case of measurement instruments, it makes it difficult to understand what the resulting measurements do and do not mean.
To address this challenge, we make a theoretical contribution to the specification of representational harms by introducing a framework grounded in speech act theory (Austin, 1962), shown at a high level in Figure 1 and described in Section 2, that conceptualizes representational harms as the perlocutionary effects, (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). This theory-grounded framework enables us to draw clearer distinctions between system behaviors and their impacts, and to provide new definitions of stereotyping, demeaning, and erasure. In Section 3, we use our framework to develop a granular taxonomy that highlights distinguishing aspects of the types of illocutionary acts—i.e., system behaviors—that cause representational harms.
@@
Figure 1: A visualization of our framework of representational harms, which we describe in detail in Section 2, instantiated with a single system behavior.
@@ 2 A Speech Act Theoretical Framework of Representational Harms
Representational harms are widely recognized among fairness-related harms caused generative language systems (e.g., Barocas et al., 2017; Crawford, 2017; Blodgett, 2021; Katzman et al., 2023). However, they are often presented as high-level concepts—specifically, stereotyping, demeaning, and erasure—with corresponding definitions that lack internal theoretical coherence and do not provide the granularity needed to develop valid measurement instruments and effective mitigations. In this section, we explain how speech act theory (Austin, 1962) is a particularly appropriate foundation for specifying representational harms.
@@
Speech act theory is a particularly appropriate foundation for specifying representational harms because of the linguistic nuance it provides. In contrast with other linguistic theories of meaning, it distinguishes between an utterance’s purpose (illocution) and real-world impacts (perlocution), while also capturing word choice and ordering (locution). As we explain, this specificity enables us to draw clearer distinctions between system behaviors and their impacts, while providing a unifying theoretical lens through which to compare different taxonomies of representational harms. Speech act theory can also be used in conjunction with other linguistic theories to provide further linguistic—specifically pragmatic, syntactic, and lexical—nuance. Consequently, it has the potential to facilitate deeper interrogations of generative language systems across a wide range of linguistic topics, including questions about the extent to which generative language systems conform to conversational norms and implicate other sociolinguistic and pragmatic frameworks. [^1] For example, we might additionally want to analyze system outputs through the lenses of Gricean maxims, adjacency pairs, common ground, implicature and presupposition, sociopragmatic variation, politeness norms, and other linguistic concepts. See the work of Birner (2013) and Huang (2014) for an overview of some of these relevant linguistic concepts.
@@ 2.1 Representational Harms as Perlocutionary Effects
Building on the argument above, we conceptualize all harms caused by generative language systems—including representational harms—as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). Specifically, representational harms occur when the perlocutionary effects of a system output include the entrenchment—i.e., the further cementing in the world—of one or more harmful222 One could argue that all social hierarchies are inherently harmful because they differentially confer power, status, privileges, resources and opportunities. That said, we intentionally include the word “harmful” to emphasize our focus on harms. For a broad range of alternative perspectives on social hierarchies and power, see Bourdieu (1984), Connell and Messerschmidt (2005), Gramsci (1971), Sidanius and Pratto (1999), Diberardino et al. (2024), and Anderson and Brown (2010). social hierarchies, where a social hierarchy is a systematic organization of individuals or social groups that differentially confers power, status, privileges, resources, and opportunities. We provide more details about individuals and social groups within harmful social hierarchies and how they relate to questions about identity formation and maintenance in Appendix B.
Harmful social hierarchies are entrenched when representations of the world that involve those hierarchies are (re-)produced—e.g., in the outputs of generative language systems. As a result, the entrenchment of one or more harmful social hierarchies constitutes a first-order perlocutionary effect because it occurs at the time of output generation. However, representations of the world that involve harmful social hierarchies can also influence individuals’ beliefs—e.g., about other individuals and social groups—as well as their psychological states—e.g., causing them to feel harmed. These real-world impacts can be either first-order or second-order perlocutionary effects because they can occur at the time of output generation or subsequent to it. We provide additional theoretical considerations relating to the entrenchment of harmful social hierarchies in Appendix B. We also explain in Appendix D how other fairness-related harms, namely allocation harms and quality-of-service harms, can be similarly conceptualized as the perlocutionary effects of particular types of illocutionary acts.
@@ 2.2 Stereotyping, Demeaning, and Erasure as Types of Illocutionary Acts
To provide new definitions of stereotyping, demeaning, and erasure, we take a top-down, or theory-first, approach. This approach differs from the ways in which existing definitions of these concepts have been developed, which include 1) adopting definitions wholesale from social psychology and other disciplines without accounting for the ways in which these discipline-specific definitions may not be a good match for generative language systems and 2) taking a bottom-up approach that uses sets of example system outputs to construct definitions. By starting with speech act theory, our top-down approach supports more cohesive understandings of stereotyping, demeaning, and erasure and the ways they manifest in language.
We conceptualize stereotyping, demeaning, and erasure as particular types of illocutionary acts, or system behaviors, whose perlocutionary effects include the entrenchment of one or more harmful social hierarchies. Collectively, these types of illocutionary acts span the five basic classes of illocutionary acts proposed by Searle (1976).
Building on and revising existing high-level definitions of stereotyping, demeaning, and erasure that relate these concepts to normative views about identity (Blodgett, 2021), we draw on speech act theory, linguistic anthropology, and sociolinguistics to provide new definitions of these concepts. Specifically, our definitions rely on the linguistic anthropological notion of evaluative lenses that can be be used to either empower or disempower individuals and social groups when using language to 1) characterize social groups, 2) position individuals within social groups, and 3) position social groups within harmful social hierarchies. In other words, these evaluative lenses can be used to either disrupt or entrench harmful social hierarchies. We consider two such evaluative lenses: similarity/difference and authenticity/inauthenticity. [^5] We adapted these evaluative lenses from a set of semiotic and social processes—called the tactics of intersubjectivity—that are identified by linguistic anthropologists as being critical to identity formation. For a review of these processes, see Bucholtz and Hall (2004) and Bucholtz and Hall (2005)’s discussions of adequation and distinction, authentication and denaturalization, and authorization and illegitimation. We include a detailed discussion of these lenses in Appendix E.
We provide our new definitions of stereotyping, demeaning, and erasure below. To illustrate these definitions, we use example utterances from ToxiGen Hartvigsen et al. (2022), a widely used, large-scale, machine-generated dataset for training classifiers to measure implicit toxicity. These examples include three utterances that target LGBTQ+ people. We present a table summarizing these examples and their corresponding types of illocutionary acts, illocutionary act patterns, evaluative lenses, and illocutionary effects in Appendix J.
@@ 2.2.1 Stereotyping
As a type of behavior, stereotyping has previously been characterized as involving stereotyping beliefs—i.e., “fixed, overgeneralized belief[s] about […] particular group[s] of people” (Cardwell, 1996). [^6] Some definitions of stereotyping focus primarily on negative sentiment. We take the view that stereotyping spans positive, negative, and neutral sentiments. This allows us to capture the variety of ways in which stereotyping manifests in language, including truly neutral cases without any particular polarized sentiment; cases that stereotype through stigmatization; and cases that stereotype through glorification of social groups, including, for example, invocations of benevolent sexism (Glick and Fiske, 1996), inspiration porn (Ladau, 2021), and the model minority myth (Walton and Truong, 2022).This definition does not specify the ways in which stereotyping manifests in language, whether all overgeneralized beliefs about particular groups of people are stereotyping, or how representations of individuals factor in. Accordingly, operationalizing this definition via measurement instruments necessarily involves relying on intuition, culture, politics, and lived experiences, leading to measurement instruments that can vary considerably (e.g., Nangia et al., 2020; Nadeem et al., 2021; Fleisig et al., 2023).
Building on this definition to add internal theoretical coherence and provide the granularity needed to develop valid measurement instruments and effective mitigations, we define a stereotyping illocutionary act to be an illocutionary act that 1) invokes both within-group similarity and between-group difference and 2) implicitly positions the referenced characteristic(s) as indicator(s) of authenticity, reinforcing the criteria by which individuals are positioned within or excluded from the targeted social group(s). These evaluative lenses are used to empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)). Specifically, a stereotyping illocutionary act essentializes the members of each targeted social group by characterizing them as being similar to one another and different from members of one or more other social groups on the basis of one or more characteristics. For example, the utterance “Gay men are not very masculine” essentializes gay men by characterizing them as being as similar to one another and, implicitly, different from straight men—on the basis of being not very masculine. The characteristic of being not very masculine is also positioned as an indicator of authenticity for gay men. We note that stereotyping illocutionary acts are typically representatives Searle (1976).
@@ 2.2.2 Demeaning
Demeaning has previously been characterized as involving “dehumanizing or offensive associations, or [behaviors that] otherwise threaten people’s sense of security or dignity” (Blodgett, 2021). This definition captures some aspects of demeaning, but stops short of specifying the ways in which it manifests in language. We build on this definition and define a demeaning illocutionary act to be an illocutionary act that disempowers one or more social groups (or one or more individuals based on their membership in those social group(s)) by invoking between-group difference and, often implicitly, within-group similarity. In contrast to a stereotyping illocutionary act, a demeaning illocutionary act does not invoke the evaluative lens of authenticity/inauthenticity. In other words, while a stereotyping illocutionary act positions the referenced characteristic(s) as indicator(s) of authenticity, a demeaning illocutionary act does not. For example, the utterance “Obviously bisexuals are messed up” disempowers bisexual people by stigmatizing them on the basis of deviance. This utterance characterizes bisexual people as being different from non-bisexual people without positioning any characteristic(s) as indicator(s) of authenticity. Unlike stereotyping illocutionary acts, which are typically representatives, demeaning illocutionary acts can be representatives, expressives, directives, or commissives Searle (1976).
@@ 2.2.3 Erasure
Finally, erasure has previously been characterized as a failure to recognize the existence of social groups, often by “foregrounding dominant understandings and […] ideologies” (Blodgett, 2021). [^7] We note that this is a contextually specific definition that aligns with established theoretical work on erasure in linguistic anthropology (see, e.g., the work of Irvine and Gal, 2000). This definition also stops short of specifying the ways in which erasure manifests in language. Moreover, erasure is especially challenging to identify because, in many cases, it involves a lack of representation. Building on this definition to add internal theoretical coherence and provide the granularity needed to develop valid measurement instruments and effective mitigations, we define an erasing illocutionary act to be an illocutionary act that invokes within-hierarchy similarity to disempower one or more social groups (or one or more individuals based on their membership those social group(s)).
An erasing illocutionary act erases differences within a harmful social hierarchy by characterizing that hierarchy as being simpler or more internally similar than it actually is, or by characterizing the members of multiple social groups as being similar to one another on the basis of one or more characteristics. In other words, an erasing illocutionary act fails to recognize socially meaningful differences within a harmful social hierarchy. For example, the utterance “There is no way that bisexuality is a real thing” presents the hierarchy of possible sexualities as simpler—e.g., comprised of a smaller set of social groups, or involving fewer characteristics—than it actually is. By invoking within-hierarchy similarity instead of between-group similarly, this definition captures an expansive understanding of erasure that highlights the role of harmful social hierarchies themselves. Erasing illocutionary acts can be representatives, expressives, directives, commissives, or declarations (Searle, 1976).
@@ 3 A Granular Taxonomy of Stereotyping, Demeaning, and Erasure
Stereotyping Describes stereotypically “All [social group] are [characteristic]” Group/individual is stereotyped
@@
Caricatures individual as a member of a social group “Those Russian trolls are always at it”
Prescribes stereotypically “[Social group] should be [characteristic]”
Proscribes stereotypically “[Social group] shouldn’t be [characteristic]”
@@
Demeaning Stigmatizes using slurs “R*tard! ” Group/individual is stigmatized
@@
Stigmatizes using general insult “Those [social group] are such jerks”
@@
Stigmatizes on basis of disorder, disease, uncleanliness, or untidiness “[Social group] are riddled with disease”
@@
Stigmatizes on basis of immorality, danger, or deviance “Being [social group] is a sin”
@@
Stigmatizes on basis of lack of social capital (popularity, likability, interest, success, or other positive quality) “[Social group] are not that successful, are they?”
@@
Advocates for stigmatizing treatment/avoidance “Avoid [social group] at all costs!”
@@
Threatens or promises stigmatizing treatment/avoidance “Better keep some distance from [social group] or you’ll end up like them”
@@
Others individuals from group as poor exemplar “They’re not a real [social group member]” Group/individual is othered
@@
Others group from (dominant) group “[Social group] are just really different from the rest of us”
@@
Others subgroup from superset group “[Social subgroup] aren’t really even [social superset group]”
@@
Advocates for treatment as other “Let’s make sure [social group] can’t get too comfortable here”
@@
Threatens or promises othering treatment “If they’re not careful, [social group] will end up being treated differently from everyone else”
@@
Objectifies (reduces to physical attribute or behavior) “[Social group] are all legs!” Group/individual is simplified
@@
Glorifies (reduces to prominent and often flattering character or personality traits) “[Social group] are so brave in the face of so much adversity”
@@
Advocates for treatment along the lines of simplified (objectifying or glorifying) features “Let’s go get us some [social group] ass”
@@
Threatens or promises treatment along the lines of simplified (objectifying or glorifying) features “We’re going to treat you [social group] like your [objectified body parts] deserve”
@@
Equates to baby/child, animal, disease or inanimate object “[Social group] are such babies” Group/individual is dehumanized
@@
Notes characteristics that match baby/child, animal, disease or inanimate object “Why are [social group] always whining?”
@@
Advocates for treatment like baby/child, animal, disease or inanimate object “[Social group] need to be disciplined”
@@
Threatens or promises treatment like baby/child, animal, disease or inanimate object “If those [social group] aren’t careful, I’m going to skewer them”
@@
Erasing Deprioritizes or questions the needs that diverge from the norm or majority “There aren’t that many [social group] here anyway, so consideration for their needs isn’t important” Group/individual needs, experiences, contributions and accountability are presented as equal
@@
Delegitimizes or denies a group’s suffering or minimizes the harm of oppression, including via denial of historical events “[Social group persecution event] wasn’t that bad for [social group] anyway”
@@
Diffuses accountability for a harm across the hierarchy, including blaming victims for their circumstances “[Social group] shouldn’t have been [characteristic] if they didn’t want that to happen to them”
@@
Advocates for non-differentiated treatment “[Social group] shouldn’t get special treatment just because they’re [characteristic]”
@@
Threatens or promises non-differentiated treatment “I guarantee we’re not going to specially accommodate [social group]”
@@
Denies justice and fair access to accommodations “No exceptions for [social group]!”
@@
Conflates individuals or social groups “Aren’t [social group 1] and [social group 2] basically the same?” Groups/individuals are homogenized and presented as indistinguishable
@@
Denies existence or fails to recognize individuals or groups “There’s no such thing as [social group]”
@@
Denies existence of individual social group members with certain characteristics “I’ve never met any [characteristic] [social group]”
Advocates for exclusion “Don’t let any [social group] in!”